# Javier Alejo, Antonio F. Galvao, Julian Martinez-Iriarte, and Gabriel Montes-Rojas

### v2

@@ A.4 Proof of Lemma 2
Since $h\equiv 1$, $V$ is identified from the first-stage in (2) as $V=D-Z^{\prime}\pi_{1}-X^{\prime}\pi_{2}$. Under Assumptions 3, 4, and 5, we have the following conditional expectation:
$E[Y|D,X,Z]$ $=D\alpha_{1}+X^{\prime}\alpha_{2}+\left(1+\sum_{s=1}^{k_{g}}(\theta_{s}^{d}D^{s}+X^{\prime s}\theta_{s}^{x})\right)\left(\sum_{j=1}^{k_{v}}\mu_{j}V^{j}\right).$ (A.5)
@@
If $h\nequiv 1$, then the first-stage in (2) is given by $D=Z^{\prime}\pi_{1}+X^{\prime}\pi_{2}+h(X,Z)V$. Consider the residual $h(X,Z)V$, and apply conditional expectations as follows:
$E[h(X,Z)V|X,Z]$ $=h(X,Z)E[V|X,Z]=0,$ (A.7)
and
$E[h(X,Z)^{2}V^{2}|X,Z]$ $=h(X,Z)^{2}E[V^{2}|X,Z]=h(X,Z)^{2}.$ (A.8)
where $E[V|X,Z]=0$ follows from Assumption 1, and $E[V^{2}|X,Z]=1$ follows from Assumption 6. Thus, it is in principle possible to identify $h(X,Z)$ in a two stage procedure. First, identify $h(X,Z)V$ as the residuals from the regression of $D$ on $Z$ and $X$, which is possible by (A.7). Second, following (A.8), run a regression of $h(X,Z)^{2}V^{2}$ on $X$ and $Z$ to obtain $h(X,Z)^{2}$. Finally, construct $V$ as
$V=\frac{D-Z^{\prime}\pi_{1}-X^{\prime}\pi_{2}}{h(X,Z)},$ (A.9)
which is possible since $h(X,Z)\neq 0$ almost surely. Once $V$ is identified, we can proceed as in the case where $h\equiv 1$. That is, $\alpha_{1}$ is identified by (9). ∎
@@ A.5 Proof of Proposition 1
Here we are going to derive the influence function of $\sqrt{n}(\widehat{\phi}-\phi)$. The procedure is detailed in section 4.1, and we reproduce it here for completeness. Following (2), the first-stage equation is $D=Z^{\prime}\pi_{1}+X^{\prime}\pi_{2}+h(X,Z;\gamma)V.$ By Assumption 1, $E[V|X,Z]=0$, so that $E[D|X,Z]=Z^{\prime}\pi_{1}+X^{\prime}\pi_{2},$ and the OLS regression of $D$ on $Z$ and $X$ yields consistent estimators of $\pi_{1}$ and $\pi_{2}$, denoted by $\widehat{\pi}_{1}$ and $\widehat{\pi}_{2}$ respectively. That is, $\widehat{\pi}_{1}$ and $\widehat{\pi}_{2}$ satisfy the empirical moment condition:
@@
The residuals of the regression are denoted by $\widetilde{V}$, and are constructed as $\widetilde{V}_{i}=D_{i}-Z_{i}^{\prime}\widehat{\pi}_{1}-X_{i}^{\prime}\widehat{\pi}_{2}.$
To estimate $\gamma$, the parameter governing the skedastic function, a model needs to be chosen. Examples are given in equations (11) and (10): $h(X,Z;\gamma)^{2}=\exp\left(\gamma_{0}+\log|Z^{\prime}|\gamma_{1}+\log|X^{\prime}|\gamma_{2}\right)$, and $h(X,Z;\gamma)^{2}=\gamma_{0}+|Z^{\prime}|\gamma_{1}+|X^{\prime}|\gamma_{2}$ respectively. By assumption, $E[V^{2}|X,Z]=1$ and $h(X,Z;\gamma)\neq 0$ almost surely, so that by equation (2)
@@
This means that a (possibly nonlinear) least squares regression of $(D-Z^{\prime}\pi_{1}-X^{\prime}\pi_{2})^{2}$ on a transformation of $Z$ and $X$ (given by either (11) or (10)) consistently estimates $\gamma$ [^14] A classic reference for nonlinear least squares regression is Jennrich (1969). A modern treatment is given in Example 5.27 of van der Vaart (1998). This is, of course, unfeasible, so we use $\tilde{V}^{2}$. To be more precise, $\widehat{\gamma}$ is obtained as
$\widehat{\gamma}=\arg\min_{a}\sum_{i=1}^{n}\left(\tilde{V}_{i}^{2}-h(X_{i},Z_{i};a)^{2}\right)^{2}$ (A.11)
@@
$\frac{1}{n}\sum_{i=1}^{n}\nabla_{\gamma}h(X_{i},Z_{i};\widehat{\gamma})\left(\tilde{V}_{i}^{2}-h(X_{i},Z_{i};\widehat{\gamma})^{2}\right)h(X_{i},Z_{i};\widehat{\gamma})=o_{p}(n^{-1/2}),$ (A.12)
@@
$+E\begin{bmatrix}-Z_{i}Z_{i}^{\prime}&-Z_{i}X_{i}^{\prime}&0\\ -X_{i}Z_{i}^{\prime}&-X_{i}X_{i}^{\prime}&0\\ -2\nabla_{\gamma}h(X_{i},Z_{i};\gamma)V_{i}Z_{i}^{\prime}h(X_{i},Z_{i};\gamma)^{2}&-2\nabla_{\gamma}h(X_{i},Z_{i};\gamma)V_{i}X_{i}^{\prime}h(X_{i},Z_{i};\gamma)^{2}&\Sigma_{\gamma}\\ \end{bmatrix}\sqrt{n}\begin{bmatrix}\widehat{\pi}_{1}-\pi_{1}\\ \widehat{\pi}_{2}-\pi_{2}\\ \widehat{\gamma}-\gamma\end{bmatrix}.$
@@
$+\begin{bmatrix}-E[Z_{i}Z_{i}^{\prime}]&-E[Z_{i}X_{i}^{\prime}]&0\\ -E[X_{i}Z_{i}^{\prime}]&-E[X_{i}X_{i}^{\prime}]&0\\ 0&0&-2E[\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\nabla_{\gamma}h(X_{i},Z_{i};\gamma)^{\prime}h(X_{i},Z_{i};\gamma)^{2}]\\ \end{bmatrix}\sqrt{n}\begin{bmatrix}\widehat{\pi}_{1}-\pi_{1}\\ \widehat{\pi}_{2}-\pi_{2}\\ \widehat{\gamma}-\gamma\end{bmatrix}.$
@@ A.6 Proof of Theorem 1
where $U:=Y-E[Y|D,W]$ and satisfies, by construction, $E[U|D,W]=0$. We follow Appendix 6A in Wooldridge (2010) closely. To implement the regression, $V$ is replaced by $\widehat{V}$. Recall that, according to (A.9), $V$ is given by
@@
$\frac{1}{n}\sum_{i=1}^{n}R_{i}(\widehat{\phi})(Y_{i}-R_{i}(\widehat{\phi})^{\prime}\widehat{\alpha})$ $=\frac{1}{n}\sum_{i=1}^{n}R_{i}(\phi)(Y_{i}-R_{i}(\phi)^{\prime}\widehat{\alpha})+\frac{1}{n}\sum_{i=1}^{n}(U_{i}J_{\phi,i}-R_{i}(\phi)\alpha^{\prime}J_{\phi,i})(\widehat{\phi}-\phi)+o_{p}(n^{-1/2}),$

### v3

@@ A.4 Proof of Lemma 2
Under Assumptions 3, 4, and 5, we have the following conditional expectation:
$E[Y|D,X,Z]$ $=D\alpha_{1}+X^{\prime}\alpha_{2}+\left(1+\sum_{s=1}^{k_{g}}(\theta_{s}^{d}D^{s}+X^{\prime s}\theta_{s}^{x})\right)\left(\sum_{j=1}^{k_{v}}\mu_{j}\tilde{V}^{j}\right).$ (A.5)
@@ A.5 Proof of Proposition 1
Here we are going to derive the influence function of $\sqrt{n}(\widehat{\phi}-\phi)$. The procedure is detailed in section 4.1, and we reproduce it here for completeness. Following Assumption 6, the first-stage equation is $D=Z^{\prime}\pi_{1}+X^{\prime}\pi_{2}+h(X,Z;\gamma)V.$ By Assumption 1, $E[V|X,Z]=0$, so that $E[D|X,Z]=Z^{\prime}\pi_{1}+X^{\prime}\pi_{2},$ and the OLS regression of $D$ on $Z$ and $X$ yields consistent estimators of $\pi_{1}$ and $\pi_{2}$, denoted by $\widehat{\pi}_{1}$ and $\widehat{\pi}_{2}$ respectively. That is, $\widehat{\pi}_{1}$ and $\widehat{\pi}_{2}$ satisfy the empirical moment condition:
@@
The residuals of the regression are denoted by $\check{V}$, and are constructed as $\check{V}_{i}=D_{i}-Z_{i}^{\prime}\widehat{\pi}_{1}-X_{i}^{\prime}\widehat{\pi}_{2}.$ To estimate $\gamma$, the parameter governing the skedastic function, a model needs to be chosen. Examples are $h(X,Z;\gamma)^{2}=\exp\left(\gamma_{0}+\log|Z^{\prime}|\gamma_{1}+\log|X^{\prime}|\gamma_{2}\right)$ and $h(X,Z;\gamma)^{2}=\gamma_{0}+|Z^{\prime}|\gamma_{1}+|X^{\prime}|\gamma_{2}$. By assumption, $E[V^{2}|X,Z]=1$ and $h(X,Z;\gamma)\neq 0$ almost surely, so that
@@
This means that a (possibly nonlinear) least squares regression of $(D-Z^{\prime}\pi_{1}-X^{\prime}\pi_{2})^{2}$ on a transformation of $Z$ and $X$ (given by either (8) or (9)) consistently estimates $\gamma$ [^8] A classic reference for nonlinear least squares regression is Jennrich (1969). A modern treatment is given in Example 5.27 of van der Vaart (1998). This is, of course, unfeasible, so we use $\check{V}^{2}$. To be more precise, $\widehat{\gamma}$ is obtained as
$\widehat{\gamma}=\arg\min_{a}\sum_{i=1}^{n}\left(\check{V}_{i}^{2}-h(X_{i},Z_{i};a)^{2}\right)^{2}$ (A.8)
@@
$\frac{1}{n}\sum_{i=1}^{n}\nabla_{\gamma}h(X_{i},Z_{i};\widehat{\gamma})\left(\check{V}_{i}^{2}-h(X_{i},Z_{i};\widehat{\gamma})^{2}\right)h(X_{i},Z_{i};\widehat{\gamma})=o_{p}(n^{-1/2}),$ (A.9)
@@
$+E\begin{bmatrix}-Z_{i}Z_{i}^{\prime}&-Z_{i}X_{i}^{\prime}&0\\ -X_{i}Z_{i}^{\prime}&-X_{i}X_{i}^{\prime}&0\\ -2\nabla_{\gamma}h(X_{i},Z_{i};\gamma)V_{i}Z_{i}^{\prime}h(X_{i},Z_{i};\gamma)^{2}&-2\nabla_{\gamma}h(X_{i},Z_{i};\gamma)V_{i}X_{i}^{\prime}h(X_{i},Z_{i};\gamma)^{2}&\Sigma_{\gamma}\\ \end{bmatrix}$
$\sqrt{n}\begin{bmatrix}\widehat{\pi}_{1}-\pi_{1}\\ \widehat{\pi}_{2}-\pi_{2}\\ \widehat{\gamma}-\gamma\end{bmatrix}.$
@@
$+\begin{bmatrix}-E[Z_{i}Z_{i}^{\prime}]&-E[Z_{i}X_{i}^{\prime}]&0\\ -E[X_{i}Z_{i}^{\prime}]&-E[X_{i}X_{i}^{\prime}]&0\\ 0&0&-2E[\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\nabla_{\gamma}h(X_{i},Z_{i};\gamma)^{\prime}h(X_{i},Z_{i};\gamma)^{2}]\\ \end{bmatrix}$
$\sqrt{n}\begin{bmatrix}\widehat{\pi}_{1}-\pi_{1}\\ \widehat{\pi}_{2}-\pi_{2}\\ \widehat{\gamma}-\gamma\end{bmatrix}.$
@@ A.6 Proof of Theorem 1
where $U:=Y-E[Y|D,W]$ and satisfies, by construction, $E[U|D,W]=0$. We follow Appendix 6A in Wooldridge (2010) closely. To implement the regression, $V$ is replaced by $\widehat{V}$. Recall that $V$ is given by
@@
$\frac{1}{n}\sum_{i=1}^{n}R_{i}(\widehat{\phi})(Y_{i}-R_{i}(\widehat{\phi})^{\prime}\widehat{\alpha})$ $=\frac{1}{n}\sum_{i=1}^{n}R_{i}(\phi)(Y_{i}-R_{i}(\phi)^{\prime}\widehat{\alpha})$
$\ \ +\frac{1}{n}\sum_{i=1}^{n}(U_{i}J_{\phi,i}-R_{i}(\phi)\alpha^{\prime}J_{\phi,i})(\widehat{\phi}-\phi)+o_{p}(n^{-1/2}),$
@@
Acknowledgements
We thank Hidehiko Ichimura, Augusto Nieto-Barthaburu, and seminar participants at Universidad de San Andrés, Universidad de Buenos Aires, and RedNIE for very helpful and constructive comments. Computer programs to replicate the numerical analyses are available from the authors. All remaining errors are our own.